# Hyperbolic Network Latent Space Model with Learnable Curvature

### v1

@@ 1 Introduction
With recent advances in science and technology, network data that contain relational information among observations are becoming more prevalent than ever. Networks emerge ubiquitously in many fields including social media (Traud et al.,, 2012), brain science (Bullmore and Sporns,, 2009), and economics (Chaney,, 2014). Statistical models are especially useful for understanding the network structure and characterizing the generative process of edges. Contrasting classical scenarios where observations are assumed to be independent, network data often display higher-order, complex dependency structures such as transitivity and modularity (Newman,, 2018).
@@
Latent space models, recognized for their flexibility in modeling complex network structures, are widely used in network analysis (Hoff et al.,, 2002; Ma et al.,, 2020; Zhang et al.,, 2022) and have demonstrated their power in capturing various features commonly observed in real-world networks (Ward and Hoff,, 2007; Ward et al.,, 2007, 2011; Friel et al.,, 2016). Under the latent space models, each node is assigned a latent position $z_{i}$, and the network edges are generated conditioning on these latent positions. One popular class of network latent space models is the distance model, first proposed in Hoff et al., (2002). It assumes that the linkage probability of two nodes $i$ and $j$ is $\sigma(d(z_{i},z_{j}))\in[0,1]$, where $d(z_{i},z_{j})$ is the distance between nodes $i$ and $j$ in the latent space. The link function $\sigma(x)$ is strictly decreasing so that nodes closer in the latent space are more likely to be connected. The model has nice properties that it naturally admits transitivity and homophily due to the distance triangular inequality: given nodes A, B, and C, if A and B, A and C are close in the latent space respectively, then B and C are close as well. In Hoff et al., (2002), the authors set $\sigma(x)=\operatorname{logistic}(\alpha-x)$ and $d(z_{i},z_{j})=|z_{i}-z_{j}|$ where $z_{i}\in\mathbb{R}$ and $\alpha$ controls the maximal linking probability. Ma et al., (2020) further studied an inner product latent space model, which in a special case can be formulated as a low-dimensional Euclidean distance model with $\sigma(x)=\operatorname{logistic}(\alpha-x^{2}/2)$ and $d(z_{i},z_{j})=\|z_{i}-z_{j}\|$.
@@
While the Euclidean space is traditionally employed as the default latent space in existing literature, its suitability for modeling network data remains an open question. In recent years, there has been accumulating evidence suggesting that hyperbolic space, a metric space characterized by constant negative curvature, may be more appropriate for network data modeling. For instance, Krioukov et al., (2010) proposed a latent space model with hyperbolic distance metric and showed that the node degree distribution naturally follows the power-law, a common characteristic of numerous real-world social and internet networks. Further, Kennedy et al., (2016) defined an analogy for curvature in discrete graphs, termed “hyperbolicity”, and used it to show that hyperbolic geometry prevails in social networks, internet networks, and collaboration networks. In the field of machine learning, researchers have proposed graph embedding models with hyperbolic geometry, reporting superior performances over their Euclidean counterparts (Nickel and Kiela,, 2017, 2018; Chami et al.,, 2019).
@@
The focus on latent space geometry in statistical modeling and inference has gained increased attention in the statistics community. Smith et al., (2019) conducted extensive simulation studies comparing the global and local characteristics of networks generated from distance models across spherical, Euclidean, and hyperbolic geometries. Their findings suggest that networks generated from hyperbolic models are more similar to networks collected from various applications than those produced from other geometries. Recently, Lubold et al., (2023) considered the problem of conducting statistical inference of the latent space curvature and proposed a model-based hypothesis testing framework. The testing procedure, motivated by the connection between the eigenvalues of the distance matrix and latent space curvature, can be used to decide whether the latent space geometry is Euclidean, hyperbolic, or spherical.
@@
However, the geometric impact of latent space curvature, as well as the optimization and theoretical analysis of the maximum-likelihood inference, remain insufficiently understood in the existing literature. Motivated by the observation that hyperbolic spaces with different curvatures are not isometric (see more details in Property 2.1), in this work we aim to integrate the geometric effect of latent space curvature into the modeling of network data. Our main contributions are as follows.
@@
First, from the modeling perspective, we propose a unified hyperbolic latent space modeling framework with a learnable curvature parameter $-K<0$, which allows us to embed networks in hyperbolic spaces with different curvatures. In particular, the model can be viewed as a generalization of the Euclidean distance model in that the hyperbolic distance becomes the Euclidean distance when $K\to 0$. Most existing hyperbolic models, e.g. Smith et al., (2019) and Nickel and Kiela, (2017), however, only considered a default curvature of $-1$ and did not recognize the geometric effect of the curvature. We further investigate the identifiability of latent space curvature and address the unanswered question about the interplay between the effect of geometry and the distribution of latent positions (Smith et al.,, 2019), demonstrating that the geometric effect of curvature is unique and cannot be substituted by the distribution of latent positions.
@@
Second, from the theoretical perspective, we establish consistency and convergence rates for the maximum-likelihood estimators of the curvature parameter and latent positions. Compared to existing literature, while Shalizi and Asta, (2017) showed that maximum-likelihood estimators of hyperbolic distance models are consistent, their result holds only when the latent space curvature is known and requires some strong asymptotic assumptions on the network linkage probabilities. Moreover, the analytic approach used in Shalizi and Asta, (2017) cannot be used to derive the consistency rates of the estimators. Our analysis adopts totally different techniques from hyperbolic geometry and low-rank matrix completion, which to the best of our knowledge establish the first consistency rate result. Notably, our theorem shows that embedding error is inevitable when the latent space curvature is misspecified, which highlights the geometric effect of curvature for all hyperbolic embedding methods beyond network latent space models.
@@
Last, from the computational perspective, we develop a manifold gradient descent algorithm as well as initialization methods to overcome the challenges in non-convex optimization of the likelihood function. We also conduct extensive simulation studies to illustrate the impact of latent space curvature and analyze a Facebook friendship network with the proposed model, demonstrating its interpretable visualization and great performance on downstream tasks.
@@
The rest of this paper is organized as follows. In Section 2 we introduce the basics of hyperbolic geometry and propose the hyperbolic latent space model for network data. Section 3 describes the estimation procedure, and the main theoretical results are presented in Section 4. To evaluate the performance and effectiveness of the proposed model, we conduct several simulation studies, the results of which are reported in Section 5, where we also propose a bootstrap-based procedure for statistical inference of the curvature parameter. In Section 6, we apply the hyperbolic latent space model to analyze a real-world Facebook friendship network, demonstrating the practical applicability of the proposed approach. Furthermore, in Section 7, we discuss potential future directions and extensions of the model. Proofs of the theoretical results are presented in the supplemental material.
@@ 2.1 Preliminary on Hyperbolic Geometry
Hyperbolic spaces are metric spaces with a constant negative sectional curvature, while Euclidean spaces have a curvature of constant 0. For hyperbolic spaces to have a well-defined notion of curvature, they must have dimensions greater than one. Intuitively, the negative curvature of hyperbolic spaces allows for more expansive room compared to Euclidean spaces, enabling them to accommodate more heterogeneous data. Specifically, the area of a hyperbolic disk grows exponentially with its radius, whereas in Euclidean space, the growth rate is quadratic. For a more detailed comparison of the basic geometric properties of hyperbolic, Euclidean, and spherical spaces, refer to Krioukov et al., (2010).
One motivating example that helps illustrate the difference between hyperbolic space and Euclidean space is the embedding of trees, which are a special type of graphical data with a hierarchical node structure. Figure 1 illustrates an example of attempting to isometrically embed a $4$-ary tree into $2$-dimensional Euclidean and hyperbolic spaces. As shown in the left plot, when the depth of the tree is only 2, the leaf nodes are quite close to each other, resulting in significant embedding distortion. On the other hand, in a two-dimensional hyperbolic space, it is possible to embed a tree with infinite depth while maintaining well-separated leaf nodes. Illustrated by the Poincaré disk model, which will be introduced in Section 2.1.1, line segments in the right plot are of the same lengths as the hyperbolic distance metric grows exponentially near the edge of the disk. In fact, it has been studied in the literature that trees with infinity depth can be nearly isometrically embedded into two-dimensional hyperbolic spaces, whereas for Euclidean space the required dimension grows exponentially with the depth of tree (Krioukov et al.,, 2010). Informally, this is because trees need exponential space for branching, and can be seen as a discrete hyperbolic space from a purely metric space perspective. Interestingly, many social networks or biological networks have local tree structures (Newman,, 2018).
Figure 1: An example of embedding tree-structured data. Left: embedded with $2$-dimensional Euclidean space; Right: embedded with $2$-dimensional Hyperbolic space.
@@ 2.1.1 Poincaré Disk Model
The Poincaré disk model typically refers to the model for a two-dimensional hyperbolic space while it can be naturally extended to a higher-dimensional space. Points of a two-dimensional hyperbolic space with curvature $-K<0$ are defined on the open unit disk of $\mathbb{R}^{2}$: $\mathbb{D}_{2}^{-K}=\{x\in\mathbb{R}^{2}~{}|~{}\|x\|_{2}<1\}.$ The corresponding distance metric of any two points $x,y\in\mathbb{D}_{2}^{-K}$, also known as the Poincaré metric, is defined as
@@
The Poincaré disk model is widely used for visualization purposes, because (1) all points lie within the open unit disk; and (2) the model is conformal, meaning angles in the Poincaré disk have identical values as if they were in the Euclidean space. However, when visualizing with the Poincaré disk model, one needs to keep in mind that the hyperbolic distance between a point and the center of the disk, grows exponentially compared with the Euclidean distance and the edge of the unit disk represents infinity. Figure 2 illustrates such property with a simple example of three points ($O$, $A$, and $B$) on the Poincaré Disk.
The model can be naturally extended to higher-dimensional hyperbolic spaces. For a $d$-dimensional hyperbolic space under Poincaré model, the set of points becomes open unit ball of $\mathbb{R}^{d}$ with the same form of distance metric:
$\mathbb{D}_{d}^{-K}=\{x\in\mathbb{R}^{d}~{}|~{}\|x\|_{2}<1\},\quad d_{\mathbb{D}_{2}^{-K}}(x,y)=\frac{1}{\sqrt{K}}\operatorname{acosh}\left(1+\frac{\|x-y\|_{2}^{2}}{(1-\|x\|_{2}^{2})(1-\|y\|^{2}_{2})}\right).$
Figure 2: Left: Geodesics of $OA$ and $OB$ on the Poincaré disk model; Right: corresponding hyperbolic and Euclidean distances for OA and OB.
@@ 2.1.2 Hyperboloid Model
To introduce the hyperboloid model for a $d$-dimensional hyperbolic space, it would be convenient to first define the Lorentzian inner product. Define the $(d+1)\times(d+1)$ diagonal matrix $\Lambda=\text{diag}(1,1,...,1,-1)$, the Lorentzian inner product is the vector product induced by $\Lambda$ on $\mathbb{R}^{d+1}$, i.e., $x^{T}\Lambda y=\sum_{i=1}^{d}x_{i}y_{i}-x_{d+1}y_{d+1}$. Under the hyperboloid model, a $d$-dimensional hyperbolic space with curvature $-K<0$ is defined on the upper hyperboloid in $\mathbb{R}^{d+1}$, denoted as $\mathbb{H}_{d}^{-K}=\{x\in\mathbb{R}^{d+1}~{}|x^{T}\Lambda x=-1,x_{d+1}>0\}$, with the corresponding distance metric to be
@@
The $d$-dimensional Poincaré ball model and hyperboloid model are equivalent, in the sense that there exists an isometry from $\mathbb{H}_{d}^{-K}$ to $\mathbb{D}_{d}^{-K}$, $F:x^{H}\in\mathbb{H}_{d}^{-K}\to x^{P}\in\mathbb{D}_{d}^{-K}$. This transformation is given by $x_{i}^{P}=F(x_{i}^{H})=x_{i}^{H}/(1+x_{d+1}^{H}),1\leq i\leq d$. Conversely, the inverse transformation $F^{-1}$ is an isometry from $\mathbb{D}_{d}^{-K}$ to $\mathbb{H}_{d}^{-K}$. We illustrate such isometry with a $2$-dimensional example in Figure 3. The isometry $F$ can be interpreted as projecting a point from the hyperboloid (in blue) onto the unit disk (in yellow), with the projection source point $(0,0,-1)$.
Figure 3: Left: triangular on the Poincaré disk model; Middle: illustration of the isometry from hyperboloid model to Poincaré disk model; Right: triangular on the hyperboloid model.
@@
Some literature (e.g., Chami et al.,, 2019) parameterize the hyperboloid model with the set of nodes and distance metric as $\mathbb{H}_{d}^{-K}=\{x\in\mathbb{R}^{d+1}\mid x^{T}\Lambda x=-1/K\}$ and $d_{\mathbb{H}^{-K}_{d}}(x,y)=\frac{1}{\sqrt{K}}\operatorname{acosh}(-Kx^{T}\Lambda y)$. The two forms are equivalent as there exists the differentiable one-to-one mapping $\phi:x\to x/\sqrt{K}$ for all points between the two forms. Similar parametrization exists for the Poincaré disk model as well.
@@
In the rest of the paper, we choose the hyperboloid model as the default model, which is favored over the Poincaré disk model for several technical reasons. From a theoretical analysis perspective, the pairwise distance matrix on the latent space can be fully characterized by products of low-rank matrices with non-linear transformations, which enables analysis techniques from the low-rank matrix completion literature to be applied. From a computational standpoint, optimizing a loss function that involves the distance metric under the hyperboloid model with a gradient-based method ends up with much simpler and stabler updating steps compared to the Poincaré disk model (Nickel and Kiela,, 2018).
Below we present some important properties of hyperbolic spaces. For more technical details, we refer readers to Bridson and Haefliger, (2013). 2.1 provides the fundamental argument for the identifiability and estimability of the curvature parameter. 2.2, similar to Euclidean self-isometries, characterizes the rotation and translation mappings in hyperbolic space.
@@
Any two $d$-dimensional hyperbolic spaces with different curvatures, denoted as $\mathbb{H}_{d}^{-K}$ and $\mathbb{H}_{d}^{-K^{\prime}}$ with $K\neq K^{\prime}$, are not isometric. In other words, there does not exist any smooth one-to-one distance-preserved mapping $\phi:\mathbb{H}_{d}^{-K}\to\mathbb{H}_{d}^{-K^{\prime}}$ such that for any $x,y\in\mathbb{H}_{d}^{-K}$, $d_{\mathbb{H}_{d}^{-K}}(x,y)=d_{\mathbb{H}_{d}^{-K^{\prime}}}(\phi(x),\phi(y))$.
@@
Every isometry of a $d$-dimensional hyperbolic space with curvature $-K<0$ to itself can be characterized with linear transformation $\mathcal{Q}$ under the hyperboloid model: $\mathcal{Q}:x\in\mathbb{H}_{d}^{-K}\to Qx\in\mathbb{H}_{d}^{-K}$, where $Q_{(d+1)\times(d+1)}$ is called a hyperboloid rotation matrix satisfying $Q\Lambda Q^{T}=\Lambda$. In the special case of $d=2$, $Q$ must be a combination of the following three basic hyperbolic rotation matrices:

### v2

@@ Abstract
Keywords: latent space model; hyperbolic space embedding; network analysis
@@ 1 Introduction
With recent advances in science and technology, network data that contain relational information among observations are becoming more prevalent than ever. Networks emerge ubiquitously in many fields including social media (Traud et al., 2012), brain science (Bullmore and Sporns, 2009), and economics (Chaney, 2014). Statistical models are especially useful for understanding the network structure and characterizing the generative process of edges. Contrasting classical scenarios where observations are assumed to be independent, network data often display higher-order, complex dependency structures such as transitivity and modularity (Newman, 2018).
@@
Latent space models, recognized for their flexibility in modeling complex network structures, are widely used in network analysis (Hoff et al., 2002; Ma et al., 2020; Zhang et al., 2022) and have demonstrated their power in capturing various features commonly observed in real-world networks (Ward and Hoff, 2007; Ward et al., 2007, 2011; Friel et al., 2016). Under the latent space models, each node is assigned a latent position $z_{i}$, and the network edges are generated conditioning on these latent positions. One popular class of network latent space models is the distance model, first proposed in Hoff et al. (2002). It assumes that the linkage probability of two nodes $i$ and $j$ is $\sigma(d(z_{i},z_{j}))\in[0,1]$, where $d(z_{i},z_{j})$ is the distance between nodes $i$ and $j$ in the latent space. The link function $\sigma(x)$ is strictly decreasing so that nodes closer in the latent space are more likely to be connected. The model has nice properties that it naturally admits transitivity and homophily due to the distance triangular inequality: given nodes A, B, and C, if A and B, A and C are close in the latent space respectively, then B and C are close as well. In Hoff et al. (2002), the authors set $\sigma(x)=\operatorname{logistic}(\alpha-x)$ and $d(z_{i},z_{j})=|z_{i}-z_{j}|$ where $z_{i}\in\mathbb{R}$ and $\alpha$ controls the maximal linking probability. Ma et al. (2020) further studied an inner product latent space model, which in a special case can be formulated as a low-dimensional Euclidean distance model with $\sigma(x)=\operatorname{logistic}(\alpha-x^{2}/2)$ and $d(z_{i},z_{j})=\|z_{i}-z_{j}\|$.
@@
While the Euclidean space is traditionally employed as the default latent space in existing literature, its suitability for modeling network data remains an open question. In recent years, there has been accumulating evidence suggesting that hyperbolic space, a metric space characterized by constant negative curvature, may be more appropriate for network data modeling. For instance, Krioukov et al. (2010) proposed a latent space model with hyperbolic distance metric and showed that the node degree distribution naturally follows the power-law, a common characteristic of numerous real-world social and internet networks. Further, Kennedy et al. (2016) defined an analogy for curvature in discrete graphs, termed “hyperbolicity”, and used it to show that hyperbolic geometry prevails in social networks, internet networks, and collaboration networks. In the field of machine learning, researchers have proposed graph embedding models with hyperbolic geometry, reporting superior performances over their Euclidean counterparts (Nickel and Kiela, 2017, 2018; Chami et al., 2019).
@@
The focus on latent space geometry in statistical modeling and inference has gained increased attention in the statistics community. Smith et al. (2019) conducted extensive simulation studies comparing the global and local characteristics of networks generated from distance models across spherical, Euclidean, and hyperbolic geometries. Their findings suggest that networks generated from hyperbolic models are more similar to networks collected from various applications than those produced from other geometries. Recently, Lubold et al. (2023) considered the problem of conducting statistical inference of the latent space curvature and proposed a model-based hypothesis testing framework. The testing procedure, motivated by the connection between the eigenvalues of the distance matrix and latent space curvature, can be used to decide whether the latent space geometry is Euclidean, hyperbolic, or spherical.
@@
However, the geometric impact of curvature on latent space modeling remains insufficiently understood in the existing literature. Motivated by the fact that Euclidean and hyperbolic spaces with different curvatures are not isometric, regardless of whether they share the same dimension (see Property 2.1 for more details), this work aims to integrate the latent space curvature into the model as a learnable parameter. Moreover, statistical inference problems related to the latent space geometry or latent positions in hyperbolic spaces have been largely underexplored in the literature. This gap motivates us to develop estimation algorithms and to establish theoretical foundations and inference procedures within the maximum likelihood framework. Our main contributions are summarized as follows.
@@
First, from the modeling perspective, we propose a unified hyperbolic latent space modeling framework with a learnable curvature parameter K<0-K<0, which allows us to embed networks in hyperbolic spaces with different curvatures. In particular, the model can be viewed as a generalization of the Euclidean distance model in that the hyperbolic distance becomes the Euclidean distance when $K\to 0$. Most existing hyperbolic models, e.g. Smith et al. (2019) and Nickel and Kiela (2017), however, only considered a default curvature of 1-1 and did not recognize the geometric effect of the curvature. We further investigate the identifiability of latent space curvature and address the unanswered question about the interplay between the effect of geometry and the distribution of latent positions (Smith et al., 2019), demonstrating that the geometric effect of curvature is unique and cannot be substituted by the distribution of latent positions.
@@
Second, from the theoretical perspective, we establish consistency and convergence rates for the maximum-likelihood estimators of the curvature parameter and latent positions. Compared to existing literature, while Shalizi and Asta (2017) showed that maximum-likelihood estimators of hyperbolic distance models are consistent, their result holds only when the latent space curvature is known and requires some strong asymptotic assumptions on the network linkage probabilities. Moreover, the analytic approach used in Shalizi and Asta (2017) cannot be used to derive the consistency rates of the estimators. Our analysis adopts totally different techniques from hyperbolic geometry and low-rank matrix completion, which to the best of our knowledge establish the first consistency rate result. Notably, our theorem shows that embedding error is inevitable when the latent space curvature is misspecified, which highlights the geometric effect of curvature for all hyperbolic embedding methods beyond network latent space models.
@@
Last, from the computational perspective, we develop a manifold gradient descent algorithm as well as initialization methods to overcome the challenges in non-convex optimization of the likelihood function. We also conduct extensive simulation studies to illustrate the impact of latent space curvature and analyze a Facebook friendship network with the proposed model, demonstrating its interpretable visualization and great performance.
@@
The rest of this paper is organized as follows. In Section 2 we introduce the basics of hyperbolic geometry and propose the hyperbolic latent space model for network data. Section 3 describes the estimation procedure, and the main theoretical results are presented in Section 4. To evaluate the performance and effectiveness of the proposed model, we conduct several simulation studies, the results of which are reported in Section 5, where we also propose a bootstrap-based procedure for statistical inference of the curvature parameter. In Section 6, we apply the hyperbolic latent space model to analyze a real-world Facebook friendship network, demonstrating the practical applicability of the proposed approach. Furthermore, in Section 7, we discuss potential future directions and extensions of the model. Proofs of the theoretical results are presented in the Supplementary Materials.
@@ 2.1 Preliminary on Hyperbolic Geometry
Hyperbolic spaces are metric spaces with a constant negative sectional curvature, while Euclidean spaces have a curvature of constant 0. For hyperbolic spaces to have a well-defined notion of curvature, they must have dimensions greater than one. Intuitively, the negative curvature of hyperbolic spaces allows for more expansive room compared to Euclidean spaces, enabling them to accommodate more heterogeneous data. Specifically, the area of a hyperbolic disk grows exponentially with its radius, whereas in Euclidean space, the growth rate is quadratic. For a more detailed comparison of the basic geometric properties of hyperbolic, Euclidean, and spherical spaces, refer to Krioukov et al. (2010).
One motivating example that helps illustrate the difference between hyperbolic space and Euclidean space is the embedding of trees, which are a special type of graphical data with a hierarchical node structure. Figure 1 illustrates an example of attempting to isometrically embed a $4$-ary tree into $2$-dimensional Euclidean and hyperbolic spaces. As shown in the left plot, when the depth of the tree is only 2, the leaf nodes are quite close to each other, resulting in significant embedding distortion. On the other hand, in a two-dimensional hyperbolic space, it is possible to embed a tree with infinite depth while maintaining well-separated leaf nodes. Illustrated by the Poincaré disk model, which will be introduced in Section 2.1.1, line segments in the right plot are of the same lengths as the hyperbolic distance metric grows exponentially near the edge of the disk. In fact, it has been studied in the literature that trees with infinity depth can be nearly isometrically embedded into two-dimensional hyperbolic spaces, whereas for Euclidean space the required dimension grows exponentially with the depth of tree (Krioukov et al., 2010). Informally, this is because trees need exponential space for branching, and can be seen as a discrete hyperbolic space from a purely metric space perspective. Interestingly, many social networks or biological networks have local tree structures (Newman, 2018).
Figure 1: An example of embedding tree-structured data. Left: embedded with 22-dimensional Euclidean space; Right: embedded with 22-dimensional Hyperbolic space.
@@ 2.1.1 Poincaré Disk Model
The Poincaré disk model typically refers to the model for a two-dimensional hyperbolic space while it can be naturally extended to a higher-dimensional space. Points of a two-dimensional hyperbolic space with curvature K<0-K<0 are defined on the open unit disk of $\mathbb{R}^{2}$: $\mathbb{D}_{2}^{-K}=\{x\in\mathbb{R}^{2}~|~\|x\|_{2}<1\}.$ The corresponding distance metric of any two points $x,y\in\mathbb{D}_{2}^{-K}$, also known as the Poincaré metric, is defined as
@@
The Poincaré disk model is widely used for visualization purposes, because (1) all points lie within the open unit disk; and (2) the model is conformal, meaning angles in the Poincaré disk have identical values as if they were in the Euclidean space. However, when visualizing with the Poincaré disk model, one needs to keep in mind that the hyperbolic distance between a point and the center of the disk, grows exponentially compared with the Euclidean distance and the edge of the unit disk represents infinity. Figure 2 illustrates such property with a simple example of three points ($O$, AA, and $B$) on the Poincaré Disk.
The model can be naturally extended to higher-dimensional hyperbolic spaces. For a $d$-dimensional hyperbolic space under Poincaré model, the set of points becomes open unit ball of $\mathbb{R}^{d}$ with the same form of distance metric as in (1).
Figure 2: Left: Geodesics of $OA$ and OBOB on the Poincaré disk model; Right: corresponding hyperbolic and Euclidean distances for OA and OB.
@@ 2.1.2 Hyperboloid Model
To introduce the hyperboloid model for a $d$-dimensional hyperbolic space, it would be convenient to first define the Lorentzian inner product. Define the $(d+1)\times(d+1)$ diagonal matrix $\Lambda=\text{diag}(1,1,...,1,-1)$, the Lorentzian inner product is the vector product induced by $\Lambda$ on $\mathbb{R}^{d+1}$, i.e., $x^{T}\Lambda y=\sum_{i=1}^{d}x_{i}y_{i}-x_{d+1}y_{d+1}$. Under the hyperboloid model, a $d$-dimensional hyperbolic space with curvature K<0-K<0 is defined on the upper hyperboloid in $\mathbb{R}^{d+1}$, denoted as $\mathbb{H}_{d}^{-K}=\{x\in\mathbb{R}^{d+1}~|x^{T}\Lambda x=-1,x_{d+1}>0\}$, with the corresponding distance metric to be
@@
The $d$-dimensional Poincaré ball model and hyperboloid model are equivalent, in the sense that there exists an isometry transformation $F$ from $\mathbb{H}_{d}^{-K}$ to $\mathbb{D}_{d}^{-K}$, $F:x^{H}\in\mathbb{H}_{d}^{-K}\to x^{P}\in\mathbb{D}_{d}^{-K}$. This transformation is given by $x_{i}^{P}=F(x_{i}^{H})=x_{i}^{H}/(1+x_{d+1}^{H}),1\leq i\leq d$. Conversely, the inverse transformation $F^{-1}$ is an isometry from $\mathbb{D}_{d}^{-K}$ to $\mathbb{H}_{d}^{-K}$. We illustrate such isometry with a $2$-dimensional example in Figure 3. The isometry $F$ can be interpreted as projecting a point from the hyperboloid (in blue) onto the unit disk (in yellow), with the projection source point (0,0,1)(0,0,-1).
Figure 3: Left: a triangle on the Poincaré disk model; Middle: isometry illustration from hyperboloid model to Poincaré disk model; Right: a triangle on the hyperboloid model.
@@
Some literature (e.g., Chami et al., 2019) parameterize the hyperboloid model with the set of nodes and distance metric as $\mathbb{H}_{d}^{-K}=\{x\in\mathbb{R}^{d+1}\mid x^{T}\Lambda x=-1/K\}$ and $d_{\mathbb{H}^{-K}_{d}}(x,y)=\frac{1}{\sqrt{K}}\operatorname{acosh}(-Kx^{T}\Lambda y)$. The two forms are equivalent as there exists the differentiable one-to-one mapping $\phi:x\to x/\sqrt{K}$ for all points between the two forms. Similar parametrization exists for the Poincaré disk model as well.
@@
In the rest of the paper, we choose the hyperboloid model as the default model, which is favored over the Poincaré disk model for several technical reasons. From a theoretical analysis perspective, the pairwise distance matrix on the latent space can be fully characterized by products of low-rank matrices with non-linear transformations, which enables analysis techniques from the low-rank matrix completion literature to be applied. From a computational standpoint, optimizing a loss function that involves the distance metric under the hyperboloid model with a gradient-based method ends up with much simpler and stabler updating steps compared to the Poincaré disk model (Nickel and Kiela, 2018).
Below we present some important properties of hyperbolic spaces. For more technical details, we refer readers to Bridson and Haefliger (2013). 2.1 provides the fundamental argument for the identifiability and estimability of the curvature parameter. 2.2, similar to Euclidean self-isometries, characterizes the rotation and translation mappings in hyperbolic space.
@@
Any two hyperbolic spaces with different curvatures, denoted as $\mathbb{H}_{d}^{-K}$ and $\mathbb{H}_{d^{\prime}}^{-K^{\prime}}$, with $K\neq K^{\prime}$, are not isometric, irrespective of whether the dimensions $d=d^{\prime}$ or $d\neq d^{\prime}$. In other words, there does not exist any smooth one-to-one distance-preserved mapping $\phi:\mathbb{H}_{d}^{-K}\to\mathbb{H}_{d}^{-K^{\prime}}$ such that for any $x,y\in\mathbb{H}_{d}^{-K}$, $d_{\mathbb{H}_{d^{\prime}}^{-K}}(x,y)=d_{\mathbb{H}_{d^{\prime}}^{-K^{\prime}}}(\phi(x),\phi(y))$.
@@
Every isometry of a $d$-dimensional hyperbolic space with curvature K<0-K<0 to itself can be characterized with linear transformation $\mathcal{Q}$ under the hyperboloid model: $\mathcal{Q}:x\in\mathbb{H}_{d}^{-K}\to Qx\in\mathbb{H}_{d}^{-K}$, where $Q_{(d+1)\times(d+1)}$ is called a hyperboloid rotation matrix satisfying $Q\Lambda Q^{T}=\Lambda$. In the special case of $d=2$, $Q$ must be a combination of the following three basic hyperbolic rotation matrices: